OpenStudy (anonymous):

Can you help me check this question? I got option D. @jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (phi):

yes, D is how you show angles marked 2 are congruent

OpenStudy (anonymous):

:DD Thank you, everyone kept saying "C" and I'm like "No..., no"..!

OpenStudy (phi):

C is true, but we want to show the *3* angles marked 3,1,2 at Q add up to 180 deg.

OpenStudy (anonymous):

Thank you ! :)

OpenStudy (aum):

I thought corresponding angles formed by parallel lines and their transversal are EQUAL not supplementary.

OpenStudy (phi):

aum, you are correct. this one *might* have a typo (or they are thinking a lot differently than I do)

OpenStudy (phi):

A is true, but I don't see a way to use it. C is true, but we would need to do a bit of explaining first... so C might be what they want.

OpenStudy (aum):

Yeah none of the choices seem good. But in the diagram they have already concluded angles marked 1 are equal and angles marked 3 are equal. So the only step that is remaining is to prove angles 1 + 2 + 3 = 180 degrees. So C comes closest, except there are three angles that are supplementary and not two angles as choice C says.

OpenStudy (aum):

I would probably pick C except the wordings should be: If two or MORE angles form a straight angle, then they are supplementary. With that we would have proved that at point Q, angles marked 1 + 2 + 3 = 180 degrees and therefore, the sum of the interior angles of the triangle marked 1 + 2 + 3 = 180 degrees.

OpenStudy (anonymous):

:/ I thought it was option "D"... >.<

OpenStudy (aum):

Option D says corresponding angles are supplementary. But corresponding angles are equal not supplementary. Read the second paragraph here: http://www.mathsisfun.com/geometry/corresponding-angles.html

OpenStudy (aum):

Corresponding angles can be supplementary only when each of the corresponding angles equal 90 degrees (a special case). But in general, corresponding angles are equal.